# Strong Linearizability without Compare&Swap: The Case of Bags

### v2

@@ 1 Introduction
Concurrent data structures are often built using linearizable implementations of objects as if they were atomic objects. Although linearizability is the usual correctness condition for concurrent algorithms, there are scenarios where composing linearizable implementations does not preserve desired properties. Namely, an algorithm that uses atomic objects may lose some of its properties when these atomic objects are replaced by linearizable implementations. Specifically, linearizability does not preserve hyperproperties (properties of sets of executions), for example, probability distributions of events for randomized programs and security properties such as noninterference [10, 6]. This can be rectified by using strongly-linearizable implementations [10]. They guarantee that the linearization of a prefix of a concurrent execution is a prefix of the linearization of the whole execution. This means that the linearization of operations in an execution prefix cannot depend on later events in the execution.
Attiya, Castaneda, and Enea [3] observed that strongly-linearizable implementations of objects typically use compare&swap objects, which can be used to solve consensus among any number of processes. They provided strongly-linearizable implementations of some objects, such as a wait-free single-writer snapshot object and a wait-free max register from a fetch&add object, and a lock-free, readable, resettable test&set object and a lock-free, readable, fetch&increment object from an infinite array of test&set objects and registers. Both fetch&add and test&set objects are less powerful than a compare&swap object.
@@
Queues and stacks have strongly-linearizable implementations from compare&swap objects [15, 21, 19]. A bag has a straightforward implementation from a stack or a queue. Therefore, a bag has a strongly-linearizable implementation from compare&swap objects. Although there is a wait-free, linearizable implementation of a stack from registers, test&set objects, and a readable fetch&increment object [1] and there is a lock-free, linearizable implementation of a queue from the same set of objects [16], neither of these implementations is strongly-linearizable. In fact, Attiya, Castañeda, and Hendler [5] proved that no lock-free, strongly-linearizable implementation of a stack or queue from such objects is possible.
@@ 2 Preliminaries
The consensus number of an object is the largest positive integer $n$ such that there is an algorithm solving consensus among $n$ processes using only instances of this object and registers. A register has consensus number 1. Another example of an object with consensus number 1 is an ABA-detecting register [2]. It supports dWrite($x$) and dRead(). When dWrite($x$) is performed, $x$ is stored in the object. When a process $p$ performs dRead(), the value stored in the object is returned, together with an output bit, which is true if and only if process $p$ has previously performed dRead() and some dWrite() has been performed since its last dRead(). We use a restricted version of an ABA-detecting register, where dWrite() does not have an input parameter and dRead() simply returns the output bit.
@@
Two operations, $op$ and $op^{\prime}$, commute if, whenever $op$ is applied to the object followed by $op^{\prime}$, the resulting value of the object is the same as when $op^{\prime}$ is applied followed by $op$. Note that the responses to $op$ and $op^{\prime}$ may be different when these operations are applied in the opposite order. The operation $op^{\prime}$ overwrites the operation $op$ if, whenever $op$ is applied to the object followed by $op^{\prime}$, the resulting value of the object is the same as when only $op^{\prime}$ is applied. Note that the response to $op^{\prime}$ may be different depending on whether $op$ was applied. An object is interfering if every two of its operations either commute or one of them overwrites the other. All interfering objects have consensus number at most 2 [12]. Registers are interfering. The test&set, resettable test&set, and fetch&increment objects and their readable versions are all interfering and have consensus number 2.
Both a stack and a queue are also examples of objects with consensus number 2 [12]. However, they are not interfering. The value of these objects is an unbounded sequence of elements, which is initially empty. A stack supports push($x$) and pop(), whereas a queue supports enqueue($x$) and dequeue(). A pop() or dequeue() that returns a value is successful. If it returns EMPTY, it is unsuccessful.
@@
There is a simple algorithm for solving consensus among 2 processes using two single-writer registers and a bag initialized with one element: To propose the value $x$, a process writes $x$ to its single-writer register and then performs a Take() operation on the bag. If the Take() returns an element, the process returns $x$. Otherwise, it reads and returns the value that the other process wrote to its single-writer register. It follows from Borowsky, Gafni, and Afek’s Initial State Lemma [7, Lemma 3.2], that consensus among 2 processes can be solved from registers and two initially empty bags. Thus, a bag has consensus number exactly 2.
@@
An implementation of an object is linearizable [14] if each of its executions has a linearization. An implementation is strongly-linearizable [10] if there is a function $f$ that maps each execution $\alpha$ to a linearization of $\alpha$ such that, for every prefix $\alpha^{\prime}$ of $\alpha$, $f(\alpha^{\prime})$ is a prefix of $f(\alpha)$.
@@ 3 Related Work
Many universal constructions (for example, from compare&swap objects and registers or from consensus objects and registers) provide strongly-linearizable, wait-free implementations of any shared object [10].
@@
Treiber [21, 19] gave a simple lock-free, strongly-linearizable implementation of a stack from a readable compare&swap object and registers. Each push or successful pop performs exactly one successful c&s, which is where the operation is linearized. Each unsuccessful pop performs exactly one read of a compare&swap object that returns NULL, which is where the operation is linearized. Michael and Scott [18] gave a lock-free, linearizable implementation of a queue from readable compare&swap objects and registers. Assuming nodes are never reused, it is possible to make a small modification to their implementation to make it strongly-linearizable [15]. Each enqueue or successful dequeue performs exactly one successful c&s to an object other than $Tail$, which is where it is linearized. Each unsuccessful dequeue performs exactly one read of a compare&swap object that returns NULL, which is where the operation is linearized.
@@
Even though there are wait-free, linearizable implementations of registers, max-registers, and single-writer snapshot objects from single-writer registers, Helmi, Higham, and Woelfel [11] proved that there are no lock-free, strongly-linearizable implementations. They also gave a wait-free, strongly-linearizable implementation of a bounded max-register from registers. Denysyuk and Woelfel [9] proved there are no wait-free, strongly-linearizable implementations of max-registers and single-writer snapshot objects from registers, although they gave lock-free strongly-linearizable implementations. Later, Ovens and Woelfel [20] gave lock-free, strongly-linearizable implementations of an ABA-detecting register and a single-writer snapshot object from a bounded number of bounded size registers.
@@
Afek, Gafni, and Morrison [1] gave a wait-free, linearizable implementation of a stack from interfering objects (specifically, a fetch&add object, registers, and test&set objects). Attiya and Enea [6] showed that this implementation is not strongly-linearizable. More generally, Attiya, Castañeda, and Hendler [5] proved that there is no lock-free, strongly-linearizable implementation of a stack or a queue shared by more than 2 processes from interfering objects.
@@
Li [16] gave a lock-free, linearizable implementation of a queue from a fetch&increment object, an infinite array of test&set objects, and an infinite array of registers. This implementation appears in Figure 1. He also gave a wait-free, linearizable implementation of a queue in which at most 2 processes may perform dequeue. It uses a fetch&increment object, an infinite array of registers, and each dequeue dynamically allocates a constant number of registers and test&set objects. David [8] gave a wait-free, linearizable implementation of a queue in which at most 1 process may enqueue. It uses registers, an infinite array of fetch&increment objects, and a two-dimensional infinite array of swap objects. It is unknown whether there is a wait-free, linearizable implementation of a queue from interfering objects in which any number of processes can enqueue and dequeue.
@@ 4 Max: a readable fetch&increment object, initialized to 1
Attiya, Castañeda, and Enea [3] claimed that Li’s lock-free implementation of a queue was also a strongly-linearizable implementation of a bag. However, their claim is false.
@@
This shows that the algorithm in Figure 1 is not a strongly-linearizable implementation of a bag for any choice of linearization points. In particular, there is a problem with the linearization points mentioned in [4]:
@@
A $b$-bounded bag shared by more than $2b$ processes has no lock-free, strongly-linearizable implementation from interfering objects. To see why, suppose $A$ is such an implementation for $n>2b$ processes. If a read operation is added to an interfering object, the result is still an interfering object. Hence, $A$ can be viewed as an implementation from readable interfering objects. Attiya, Castaneda, and Enea [3] defined a $b$-ordering object. A $b$-bounded bag is an example of a $b$-ordering object. Applying Lemma 5.2 from their paper shows that $b$-set agreement among $n$ processes can be solved using $A$. Since all interfering objects have consensus number at most 2, this implies that $b$-set agreement among $n>2b$ processes can be solved using objects of consensus number at most 2. Herlihy and Rajsbaum [13] proved that this is impossible.
@@ 4.1 A linearizable, but not strongly-linearizable, queue
It is worth noting that the algorithm in Figure 2 also implements a linearizable queue. The proof of this fact is very similar to the proof used by Li [16]. However, it is not a strongly-linearizable queue, as we next prove.
@@ 14 Allocated: a single-writer register, initialized to 1, which can only be written to by the producer
Our first bounded bag algorithm from interfering objects is a wait-free, linearizable implementation of a bag that can contain at most one element. It is shared by $n$ processes, $P_{1},\ldots,P_{n}$, called consumers, that can perform Take() and a single process, called the producer, that can perform Insert($x$). To keep the space bounded, the producer reuses locations in Items to store new elements. The producer announces the most recent location it has allocated by writing it to the shared register Allocated. There is a resettable test&set object, instead of a test&set object, associated with each location in Items. An array, Hazards, of hazard pointers [17] is used to prevent multiple consumers from returning the same element, and to ensure that each element is consumed before being overwritten by a new element. Each consumer announces a location it is about to access and the producer avoids reusing the announced locations. Specifically, the producer avoids resetting the corresponding test&set objects and writing new elements to the corresponding locations in Items.
@@ 5.1 Linearizability
Proof 1
All components of Items are initialized to $\bot$. The first successful Insert writes an element to Items[Allocated] on LABEL:WF-1-bounded:insert:write_item. Any subsequent successful Insert writes $\bot$ to Items[Allocated] on LABEL:WF-1-bounded:insert:clear_item. Note that the producer does not change the value of Allocated between LABEL:WF-1-bounded:insert:write_item and LABEL:WF-1-bounded:insert:clear_item, so, immediately after LABEL:WF-1-bounded:insert:clear_item, all components of Items are $\bot$. Later, on LABEL:WF-1-bounded:insert:write_item, the producer writes an element to Items[Allocated].
@@
Proof 2
The producer adds the location m = $k$ to used on LABEL:WF-1-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:WF-1-bounded:insert:clear_item. Thus, Items[k] = $\bot$ when $k$ is added to used. Suppose that an element is subsequently written to Items[k]. This occurs on LABEL:WF-1-bounded:insert:write_item, with m = $k$. The value of m was chosen on LABEL:WF-1-bounded:insert:choose_index to be a value that is not in hazardous. Thus, if $k$ was in used, it was removed from used on LABEL:WF-1-bounded:insert:narrow_used_down, before an element was written to Items[m].
@@
Right after $D_{2}$, a t&s of TS[a] succeeds, so TS[a] = 0 at $D_{2}$. Assume, for sake of contradiction, that TS[a] = 1 at some point between $D_{0}$ and $D_{2}$. Then, after this point, but before $tk$ performed its successful t&s on TS[a], the producer must have reset TS[a] on LABEL:WF-1-bounded:insert:reset_TS. By Lemma 1, Items[a] = $\bot$ when the reset was performed. Thus, between $D_{1}$ and $D_{2}$, the producer must have written $\bot$ to Items[a] on LABEL:WF-1-bounded:insert:clear_item and then performed the reset. Between these two steps, the producer collects Hazards and reads $a$ from Hazards[i]. But, by LABEL:WF-1-bounded:insert:reset_loop, the producer does not reset TS[a] on LABEL:WF-1-bounded:insert:reset_TS. This is a contradiction. Thus TS[a] = 0 between $D_{0}$ and $D_{2}$.
After writing $x$ to Items[a] just before $D_{0}$, the producer next writes to Items[a] on LABEL:WF-1-bounded:insert:clear_item. Between these two writes, the producer must read 1 from TS[a] on LABEL:WF-1-bounded:insert:check_TS[m]. But TS[a] = 0 between $D_{0}$ and $D_{2}$. Thus, Items[a] = $x$ between $D_{0}$ and $D_{2}$.
@@
So in both cases $a$ was added to used prior to $e$. By Lemma 2, Items[a] = $\bot$ while $a$ was in used. Immediately after step $e$, Items[a] $\neq\bot$, so $a$ is not in used. Prior to being removed from used on LABEL:WF-1-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:WF-1-bounded:insert:reset_TS by some step $e^{\prime}$.
By Lemma 1, Items[a] = $\bot$ at $e^{\prime}$. Between $e^{\prime}$ and $e$, there was no write to Items[a] on LABEL:WF-1-bounded:insert:write_item, so Items[a] = $\bot$. When a successful t&s on TS[a] is performed, Corollary 1 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s on TS[a] was performed between $e^{\prime}$ and $e$. This implies that TS[a] = 0 when $e$ was performed.
@@
Otherwise, $tk$ reads $\bot$ from Items[a]. But when $tk$ executes step $e$, Items[a] $\neq$ $\bot$. By Lemma 1, at that point the producer must be strictly between LABEL:WF-1-bounded:insert:write_item and the following execution of LABEL:WF-1-bounded:insert:clear_item. Let $e^{\prime}$ be the step in which the producer entered this interval, namely, the last step of the execution in which it writes an element to Items on LABEL:WF-1-bounded:insert:write_item before $e$. As the producer does not reach LABEL:WF-1-bounded:insert:clear_item between $e^{\prime}$ and $e$, it specifically does not reach LABEL:WF-1-bounded:insert:write_to_Allocated (in which it modifies Allocated) during that interval. Hence, it wrote to location $a$ at $e^{\prime}$. By Lemma 4, TS[a] = 0 at $e^{\prime}$. TS[a] = 0 at $e$ by assumption. It must remain 0 between $e^{\prime}$ and $e$, because if it had been set to 1, the producer would have had to reset it before $e$, but the producer does not reach LABEL:WF-1-bounded:insert:reset_TS during this interval. Since $tk$ reads $\bot$ from Items[a] after $e$, it indicates that the producer previously wrote $\bot$ to Items[a] on LABEL:WF-1-bounded:insert:clear_item. Before this write (and after $e^{\prime}$) the producer read 1 from TS[a] on LABEL:WF-1-bounded:insert:check_TS[m]. This implies that a successful t&s of TS[a] was executed after $e^{\prime}$. We have shown it could not occur between $e^{\prime}$ and $e$, so it occurred between $e$ and the execution of LABEL:WF-1-bounded:take:read_item by $tk$.
The linearization point of each operation occurs at one of its own steps, except for an unsuccessful Take operation $tk$ in case Items[a] $\neq$ $\bot$ and TS[a] = 0 when it reads Allocated. In this case, it is linearized after this read and before it completes, by Lemma 5. Thus, in all cases, the linearization point of an operation is within its operation interval.
@@
Let $D_{1}$ be the configuration right before this read and let $D_{2}$ be the configuration at the end of $\sigma$. Let $D_{0}$ be the configuration right after the last write to Items[a] on LABEL:WF-1-bounded:insert:write_item that was performed by the producer before $D_{1}$ and let insins be the Insert operation that performed this write. By Lemma 3, TS[a] = 0 and Items[a] = $x$ between $D_{0}$ and $D_{2}$. Therefore, by Lemma 1, no write to any cell of Items on LABEL:WF-1-bounded:insert:write_item is performed between $D_{0}$ and $D_{2}$, which means insins performed the last write of an element to Items in $\sigma$.
@@
Since insins performed LABEL:WF-1-bounded:insert:write_item right before $D_{0}$, it was completed in $\sigma$. Hence insins appears in $S$. Immediately after it appears, the bag contains only $x$. Between $D_{0}$ and $D_{2}$, TS[a] = 0, so no successful t&s is performed on TS[a]. Also, Items[a] $\neq\bot$, so, by Lemma 1, Items[a’] $\bot$ for all $a^{\prime}\neq a$. By Corollary 1, Items[a’] $\neq\bot$ when a successful t&s is performed on TS[a’]. Thus, no successful t&s is performed on any location of TS between $D_{0}$ and $D_{2}$. Hence, no Take operation that removes $x$ is linearized between $D_{0}$ and $D_{2}$. Since insins is linearized right before $D_{0}$, there are no successful Take operations after insins in $S$. Therefore, at the end of $S$, the bag still contains $x$. Since $S$ is a linearization, it follows that $B=\{x\}$.

### v3

@@ Abstract
The conference version of this paper is available at [ellen2025strong].
@@ 1 Introduction
Concurrent data structures are often built using linearizable implementations of objects as if they were atomic objects. Although linearizability is the usual correctness condition for concurrent algorithms, there are scenarios where composing linearizable implementations does not preserve desired properties. Namely, an algorithm that uses atomic objects may lose some of its properties when these atomic objects are replaced by linearizable implementations. Specifically, linearizability does not preserve hyperproperties (properties of sets of executions), for example, probability distributions of events for randomized programs and security properties such as noninterference [golab2011linearizable, attiya2019putting]. This can be rectified by using strongly-linearizable implementations [golab2011linearizable]. They guarantee that the linearization of a prefix of a concurrent execution is a prefix of the linearization of the whole execution. This means that the linearization of operations in an execution prefix cannot depend on later events in the execution.
Attiya, Castaneda, and Enea [attiya2024strong] observed that strongly-linearizable implementations of objects typically use compare&swap objects, which can be used to solve consensus among any number of processes. They provided strongly-linearizable implementations of some objects, such as a wait-free single-writer snapshot object and a wait-free max register from a fetch&add object, and a lock-free, readable, resettable test&set object and a lock-free, readable, fetch&increment object from an infinite array of test&set objects and registers. Both fetch&add and test&set objects are less powerful than a compare&swap object.
@@
Queues and stacks have strongly-linearizable implementations from compare&swap objects [hwang2022strongly, treiber1986systems, michael1998nonblocking]. A bag has a straightforward implementation from a stack or a queue. Therefore, a bag has a strongly-linearizable implementation from compare&swap objects. Although there is a wait-free, linearizable implementation of a stack from registers, test&set objects, and a readable fetch&increment object [afek2007common2] and there is a lock-free, linearizable implementation of a queue from the same set of objects [Li01], neither of these implementations is strongly-linearizable. In fact, Attiya, Castañeda, and Hendler [ACH18] proved that no lock-free, strongly-linearizable implementation of a stack or queue from such objects is possible.
@@ 2 Preliminaries
The consensus number of an object is the largest positive integer $n$ such that there is an algorithm solving consensus among $n$ processes using only instances of this object and registers. A register has consensus number 1. Another example of an object with consensus number 1 is an ABA-detecting register [aghazadeh2015time]. It supports dWrite($x$) and dRead(). When dWrite($x$) is performed, $x$ is stored in the object. When a process $p$ performs dRead(), the value stored in the object is returned, together with an output bit, which is true if and only if process $p$ has previously performed dRead() and some dWrite() has been performed since its last dRead(). We use a restricted version of an ABA-detecting register, where dWrite() does not have an input parameter and dRead() simply returns the output bit.
@@
Two operations, $op$ and $op^{\prime}$, commute if, whenever $op$ is applied to the object followed by $op^{\prime}$, the resulting value of the object is the same as when $op^{\prime}$ is applied followed by $op$. Note that the responses to $op$ and $op^{\prime}$ may be different when these operations are applied in the opposite order. The operation $op^{\prime}$ overwrites the operation $op$ if, whenever $op$ is applied to the object followed by $op^{\prime}$, the resulting value of the object is the same as when only $op^{\prime}$ is applied. Note that the response to $op^{\prime}$ may be different depending on whether $op$ was applied. An object is interfering if every two of its operations either commute or one of them overwrites the other. All interfering objects have consensus number at most 2 [herlihy1991wait]. Registers are interfering. The test&set, resettable test&set, and fetch&increment objects and their readable versions are all interfering and have consensus number 2.
Both a stack and a queue are also examples of objects with consensus number 2 [herlihy1991wait]. However, they are not interfering. The value of these objects is an unbounded sequence of elements, which is initially empty. A stack supports push($x$) and pop(), whereas a queue supports enqueue($x$) and dequeue(). A pop() or dequeue() that returns a value is successful. If it returns EMPTY, it is unsuccessful.
@@
There is a simple algorithm for solving consensus among 2 processes using two single-writer registers and a bag initialized with one element: To propose the value $x$, a process writes $x$ to its single-writer register and then performs a Take() operation on the bag. If the Take() returns an element, the process returns $x$. Otherwise, it reads and returns the value that the other process wrote to its single-writer register. It follows from Borowsky, Gafni, and Afek’s Initial State Lemma [borowsky1994consensus, Lemma 3.2], that consensus among 2 processes can be solved from registers and two initially empty bags. Thus, a bag has consensus number exactly 2.
@@
An implementation of an object is linearizable [HerlihyWing90] if each of its executions has a linearization. An implementation is strongly-linearizable [golab2011linearizable] if there is a function $f$ that maps each execution $\alpha$ to a linearization of $\alpha$ such that, for every prefix $\alpha^{\prime}$ of $\alpha$, $f(\alpha^{\prime})$ is a prefix of $f(\alpha)$.
@@ 3 Related Work
Many universal constructions (for example, from compare&swap objects and registers or from consensus objects and registers) provide strongly-linearizable, wait-free implementations of any shared object [golab2011linearizable].
@@
Treiber [treiber1986systems, michael1998nonblocking] gave a simple lock-free, strongly-linearizable implementation of a stack from a readable compare&swap object and registers. Each push or successful pop performs exactly one successful c&s, which is where the operation is linearized. Each unsuccessful pop performs exactly one read of a compare&swap object that returns NULL, which is where the operation is linearized.
Michael and Scott [michael1996simple] gave a lock-free, linearizable implementation of a queue from readable compare&swap objects and registers. Assuming nodes are never reused, it is possible to make a small modification to their implementation to make it strongly-linearizable [hwang2022strongly]. Each enqueue or successful dequeue performs exactly one successful c&s to an object other than $Tail$, which is where it is linearized. Each unsuccessful dequeue performs exactly one read of a compare&swap object that returns NULL, which is where the operation is linearized.
@@
Even though there are wait-free, linearizable implementations of registers, max-registers, and single-writer snapshot objects from single-writer registers, Helmi, Higham, and Woelfel [HHW12] proved that there are no lock-free, strongly-linearizable implementations. They also gave a wait-free, strongly-linearizable implementation of a bounded max-register from registers. Denysyuk and Woelfel [DW2015] proved there are no wait-free, strongly-linearizable implementations of max-registers and single-writer snapshot objects from registers, although they gave lock-free strongly-linearizable implementations. Later, Ovens and Woelfel [ovens2019strongly] gave lock-free, strongly-linearizable implementations of an ABA-detecting register and a single-writer snapshot object from a bounded number of bounded size registers.
@@
Afek, Gafni, and Morrison [afek2007common2] gave a wait-free, linearizable implementation of a stack from interfering objects (specifically, a fetch&add object, registers, and test&set objects). Attiya and Enea [attiya2019putting] showed that this implementation is not strongly-linearizable. More generally, Attiya, Castañeda, and Hendler [ACH18] proved that there is no lock-free, strongly-linearizable implementation of a stack or a queue shared by more than 2 processes from interfering objects.
@@
Li [Li01] gave a lock-free, linearizable implementation of a queue from a fetch&increment object, an infinite array of test&set objects, and an infinite array of registers. This implementation appears in Figure 1. He also gave a wait-free, linearizable implementation of a queue in which at most 2 processes may perform dequeue. It uses a fetch&increment object, an infinite array of registers, and each dequeue dynamically allocates a constant number of registers and test&set objects. David [D04] gave a wait-free, linearizable implementation of a queue in which at most 1 process may enqueue. It uses registers, an infinite array of fetch&increment objects, and a two-dimensional infinite array of swap objects. It is unknown whether there is a wait-free, linearizable implementation of a queue from interfering objects in which any number of processes can enqueue and dequeue.
@@ 4 Max: a readable fetch&increment object, initialized to 1
Attiya, Castañeda, and Enea [attiya2024strong] claimed that Li’s lock-free implementation of a queue was also a strongly-linearizable implementation of a bag. However, their claim is false.
@@
This shows that the algorithm in Figure 1 is not a strongly-linearizable implementation of a bag for any choice of linearization points. In particular, there is a problem with the linearization points mentioned in [attiya2024strongArxiv]:
@@
A $b$-bounded bag shared by more than $2b$ processes has no lock-free, strongly-linearizable implementation from interfering objects. To see why, suppose $A$ is such an implementation for $n>2b$ processes. If a read operation is added to an interfering object, the result is still an interfering object. Hence, $A$ can be viewed as an implementation from readable interfering objects. Attiya, Castaneda, and Enea [attiya2024strong] defined a $b$-ordering object. A $b$-bounded bag is an example of a $b$-ordering object. Applying Lemma 5.2 from their paper shows that $b$-set agreement among $n$ processes can be solved using $A$. Since all interfering objects have consensus number at most 2, this implies that $b$-set agreement among $n>2b$ processes can be solved using objects of consensus number at most 2. Herlihy and Rajsbaum [HR94] proved that this is impossible.
@@ 4.1 A linearizable, but not strongly-linearizable, queue
It is worth noting that the algorithm in Figure 2, which implements a strongly-linearizable bag, also implements a linearizable queue. The proof of this fact is very similar to the proof used by Li [Li01]. However, it is not a strongly-linearizable queue, as we next prove.
@@ 14 Allocated: a single-writer register, initialized to 1, which can only be written to by the producer
Our first bounded bag algorithm from interfering objects is a wait-free, linearizable implementation of a bag that can contain at most one element. It is shared by $n$ processes, $P_{1},\ldots,P_{n}$, called consumers, that can perform Take() and a single process, called the producer, that can perform Insert($x$). To keep the space bounded, the producer reuses locations in Items to store new elements. The producer announces the most recent location it has allocated by writing it to the shared register Allocated. There is a resettable test&set object, instead of a test&set object, associated with each location in Items. An array, Hazards, of hazard pointers [michael2004hazard] is used to prevent multiple consumers from returning the same element, and to ensure that each element is consumed before being overwritten by a new element. Each consumer announces a location it is about to access and the producer avoids reusing the announced locations. Specifically, the producer avoids resetting the corresponding test&set objects and writing new elements to the corresponding locations in Items.
@@ 5.1 Linearizability
Proof.
All components of Items are initialized to $\bot$. The first successful Insert writes an element to Items[Allocated] on LABEL:WF-1-bounded:insert:write_item. Any subsequent successful Insert writes $\bot$ to Items[Allocated] on LABEL:WF-1-bounded:insert:clear_item. Note that the producer does not change the value of Allocated between LABEL:WF-1-bounded:insert:write_item and LABEL:WF-1-bounded:insert:clear_item, so, immediately after LABEL:WF-1-bounded:insert:clear_item, all components of Items are $\bot$. Later, on LABEL:WF-1-bounded:insert:write_item, the producer writes an element to Items[Allocated]. ∎
@@
Proof.
The producer adds the location m = $k$ to used on LABEL:WF-1-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:WF-1-bounded:insert:clear_item. Thus, Items[k] = $\bot$ when $k$ is added to used. Suppose that an element is subsequently written to Items[k]. This occurs on LABEL:WF-1-bounded:insert:write_item, with m = $k$. The value of m was chosen on LABEL:WF-1-bounded:insert:choose_index to be a value that is not in hazardous. Thus, if $k$ was in used, it was removed from used on LABEL:WF-1-bounded:insert:narrow_used_down, before an element was written to Items[m]. ∎
@@
Right after $D_{2}$, a t&s of TS[a] succeeds, so TS[a] = 0 at $D_{2}$. Assume, for sake of contradiction, that TS[a] = 1 at some point between $D_{0}$ and $D_{2}$. Then, after this point, but before $tk$ performed its successful t&s on TS[a], the producer must have reset TS[a] on LABEL:WF-1-bounded:insert:reset_TS. By Lemma 5.1, Items[a] = $\bot$ when the reset was performed. Thus, between $D_{1}$ and $D_{2}$, the producer must have written $\bot$ to Items[a] on LABEL:WF-1-bounded:insert:clear_item and then performed the reset. Between these two steps, the producer collects Hazards and reads $a$ from Hazards[i]. But, by LABEL:WF-1-bounded:insert:reset_loop, the producer does not reset TS[a] on LABEL:WF-1-bounded:insert:reset_TS. This is a contradiction. Thus TS[a] = 0 between $D_{0}$ and $D_{2}$.
After writing $x$ to Items[a] just before $D_{0}$, the producer next writes to Items[a] on LABEL:WF-1-bounded:insert:clear_item. Between these two writes, the producer must read 1 from TS[a] on LABEL:WF-1-bounded:insert:check_TS[m]. But TS[a] = 0 between $D_{0}$ and $D_{2}$. Thus, Items[a] = $x$ between $D_{0}$ and $D_{2}$. ∎
@@
So in both cases $a$ was added to used prior to $e$. By Lemma 5.2, Items[a] = $\bot$ while $a$ was in used. Immediately after step $e$, Items[a] $\neq\bot$, so $a$ is not in used. Prior to being removed from used on LABEL:WF-1-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:WF-1-bounded:insert:reset_TS by some step $e^{\prime}$.
By Lemma 5.1, Items[a] = $\bot$ at $e^{\prime}$. Between $e^{\prime}$ and $e$, there was no write to Items[a] on LABEL:WF-1-bounded:insert:write_item, so Items[a] = $\bot$. When a successful t&s on TS[a] is performed, Corollary 5.4 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s on TS[a] was performed between $e^{\prime}$ and $e$. This implies that TS[a] = 0 when $e$ was performed. ∎
@@
Otherwise, $tk$ reads $\bot$ from Items[a]. But when $tk$ executes step $e$, Items[a] $\neq$ $\bot$. By Lemma 5.1, at that point the producer must be strictly between LABEL:WF-1-bounded:insert:write_item and the following execution of LABEL:WF-1-bounded:insert:clear_item. Let $e^{\prime}$ be the step in which the producer entered this interval, namely, the last step of the execution in which it writes an element to Items on LABEL:WF-1-bounded:insert:write_item before $e$. As the producer does not reach LABEL:WF-1-bounded:insert:clear_item between $e^{\prime}$ and $e$, it specifically does not reach LABEL:WF-1-bounded:insert:write_to_Allocated (in which it modifies Allocated) during that interval. Hence, it wrote to location $a$ at $e^{\prime}$. By Lemma 5.5, TS[a] = 0 at $e^{\prime}$. TS[a] = 0 at $e$ by assumption. It must remain 0 between $e^{\prime}$ and $e$, because if it had been set to 1, the producer would have had to reset it before $e$, but the producer does not reach LABEL:WF-1-bounded:insert:reset_TS during this interval. Since $tk$ reads $\bot$ from Items[a] after $e$, it indicates that the producer previously wrote $\bot$ to Items[a] on LABEL:WF-1-bounded:insert:clear_item. Before this write (and after $e^{\prime}$) the producer read 1 from TS[a] on LABEL:WF-1-bounded:insert:check_TS[m]. This implies that a successful t&s of TS[a] was executed after $e^{\prime}$. We have shown it could not occur between $e^{\prime}$ and $e$, so it occurred between $e$ and the execution of LABEL:WF-1-bounded:take:read_item by $tk$. ∎
The linearization point of each operation occurs at one of its own steps, except for an unsuccessful Take operation $tk$ in case Items[a] $\neq$ $\bot$ and TS[a] = 0 when it reads Allocated. In this case, it is linearized after this read and before it completes, by Lemma 5.6. Thus, in all cases, the linearization point of an operation is within its operation interval.
@@
Let $D_{1}$ be the configuration right before this read and let $D_{2}$ be the configuration at the end of $\sigma$. Let $D_{0}$ be the configuration right after the last write to Items[a] on LABEL:WF-1-bounded:insert:write_item that was performed by the producer before $D_{1}$ and let insins be the Insert operation that performed this write. By Lemma 5.3, TS[a] = 0 and Items[a] = $x$ between $D_{0}$ and $D_{2}$. Therefore, by Lemma 5.1, no write to any cell of Items on LABEL:WF-1-bounded:insert:write_item is performed between $D_{0}$ and $D_{2}$, which means insins performed the last write of an element to Items in $\sigma$.
@@
Since insins performed LABEL:WF-1-bounded:insert:write_item right before $D_{0}$, it was completed in $\sigma$. Hence insins appears in $S$. Immediately after it appears, the bag contains only $x$. Between $D_{0}$ and $D_{2}$, TS[a] = 0, so no successful t&s is performed on TS[a]. Also, Items[a] $\neq\bot$, so, by Lemma 5.1, Items[a’] $\bot$ for all $a^{\prime}\neq a$. By Corollary 5.4, Items[a’] $\neq\bot$ when a successful t&s is performed on TS[a’]. Thus, no successful t&s is performed on any location of TS between $D_{0}$ and $D_{2}$. Hence, no Take operation that removes $x$ is linearized between $D_{0}$ and $D_{2}$. Since insins is linearized right before $D_{0}$, there are no successful Take operations after insins in $S$. Therefore, at the end of $S$, the bag still contains $x$. Since $S$ is a linearization, it follows that $B=\{x\}$.